OpenStudy (lazyboy):

1/3, 1, 5/3, 7/3, … Question 9 answers 3, 10/3, 4 11/3, 13/3, 5 3, 11/3, 13/3 10/3, 4, 14/3

terenzreignz (terenzreignz):

something bothering you? Maybe if you write 1 as 3/3, it'll be clearer :D

OpenStudy (lazyboy):

? :D

terenzreignz (terenzreignz):

Well, if you write 1 as 3/3, a pattern emerges \[\huge \frac13\frac33\frac53\frac73\] Clear now?

OpenStudy (lazyboy):

hmm., so whats the answer? brb

terenzreignz (terenzreignz):

You just keep adding 2/3

OpenStudy (lazyboy):

I'm back

terenzreignz (terenzreignz):

And you have my observation :D

OpenStudy (lazyboy):

man u don't give me direct answers..:D ]a.b.c or d?

terenzreignz (terenzreignz):

It's actually heavily discouraged by the Code of Conduct... plus, it's more rewarding if you figure them out yourself... I already told you how to do it, namely, just keep adding 2/3 ;)

OpenStudy (lazyboy):

..! :D

terenzreignz (terenzreignz):

So... what's \[\large \frac73 + \frac23\]?

OpenStudy (lazyboy):

so A

terenzreignz (terenzreignz):

Maybe :) Why A?

OpenStudy (lazyboy):

cause I used my calculator :D

terenzreignz (terenzreignz):

Yeah, but A isn't the only choice that starts with 3.

OpenStudy (lazyboy):

oh it's cd duuuh :D

OpenStudy (lazyboy):

C

terenzreignz (terenzreignz):

Continuing... \[\huge \frac13\frac33\frac53\frac73\frac93\frac{11}3\frac{13}3\frac{15}3\]

OpenStudy (lazyboy):

yeah so C